### v3

# Integrated Sensing and Communications Over the Years: An Evolution Perspective

Abstract
Integrated sensing and communications (ISAC) enables efficient spectrum utilization and reduces hardware costs for beyond 5G (B5G) and 6G networks, facilitating intelligent applications that require both high-performance communication and precise sensing capabilities. This survey provides a comprehensive review of the evolution of ISAC over the years. We examine the expansion of spectrum across radio-frequency (RF) and optical ISAC, highlighting the role of advanced technologies, along with key challenges and synergies. We further discuss the advancements in network architecture from single-cell to multi-cell systems, emphasizing the integration of collaborative sensing and interference mitigation strategies. Moreover, we analyze the progress from single-modal to multi-modal sensing, with a focus on the integration of edge intelligence to enable real-time data processing, reduce latency, and enhance decision-making. Finally, we extensively review standardization efforts by 3GPP, IEEE, and ITU, examining the transition of ISAC-related technologies and their implications for the deployment of 6G networks.
Index Terms:
Integrated Sensing and Communications (ISAC), waveform design, optical ISAC, network architecture, edge perception, security and privacy, 6G.
I Introduction
I-A Background and Motivation
Wireless communication networks are widely regarded as foundational and strategic industries, serving as the backbone for emerging sectors such as the industrial Internet of Things (IIoT) and intelligent manufacturing [cui2025integratedsensingcommunicationmultifunctional]. Their continued evolution is considered a critical element in global science and technology agendas. Since the advent of radio systems in 1934, the independent development of communication and radar systems has introduced several challenges, including inefficient and competitive use of wireless resources [10255711], duplication of infrastructure and hardware [zheng2019radar], and limited cross-system interoperability [8828023]. Given that electromagnetic signals inherently support both environmental sensing and data transmission, maintaining separate system designs has become increasingly redundant.
TABLE I: Comparison of Survey Focus Areas and Application Scenarios
Reference Focus of Discussion Application Scenarios
RF to
Optical Single-Cell to
Multi-Cell Single-Modal
to Multi-Modal Security and Privacy Standardization Process
[9606831] $\times$ $\triangle$ $\times$ $\times$ $\times$ IoT applications
[9737357] $\times$ $\triangle$ $\times$ $\times$ $\triangle$ Dual-functional wireless networks in 6G
[9585321] $\times$ $\triangle$ $\times$ $\times$ Perceptive mobile network structure, signal processing
[wei2024deep] $\times$ $\times$ $\times$ $\times$ Resource, node, and infrastructure-level cooperation
[10418473] $\times$ $\triangle$ $\times$ $\times$ $\triangle$ Vehicular networks, 6G applications
[9296833] $\times$ $\times$ $\times$ $\times$ Perceptive mobile networks
[10049304] $\times$ $\triangle$ $\times$ $\times$ Multi-view sensing architectures and designs
[10049817] $\times$ $\triangle$ $\times$ $\times$ $\times$ ISAC hardware testbeds and performance evaluation
[8828030] $\times$ $\triangle$ $\times$ $\times$ $\times$ Signal processing for ISAC
[10.1145/3310194] $\times$ $\triangle$ $\times$ $\times$ $\times$ Wi-Fi / CSI-based sensing
[10188491] $\times$ $\triangle$ $\times$ $\times$ $\triangle$ Historical development of ISAC
[9705498] $\times$ $\times$ $\times$ $\times$ Theoretical performance bounds
[8999605] $\triangle$ $\times$ $\times$ $\times$ $\times$ Spectrum and hardware sharing
[9540344] $\times$ $\triangle$ $\times$ $\times$ $\times$ Waveform design and MIMO processing
[10004900] $\times$ $\triangle$ $\times$ $\times$ $\times$ UAV network structure
[10770016] $\times$ $\triangle$ $\triangle$ $\times$ $\times$ Multi-carrier waveform design and learning techniques
[10726912] $\times$ $\times$ $\times$ $\times$ Multi-cell ISAC networks, interference management
[11015430] $\times$ $\triangle$ $\triangle$ $\triangle$ AI/ML-assisted ISAC architecture and secure implementation
[10536135] $\triangle$ $\triangle$ $\times$ $\times$ $\triangle$ Visionary techniques for 6G ISAC
[10781415] $\times$ $\times$ $\times$ $\times$ $\times$ RIS-NOMA-based ISAC techniques
[10608156] $\triangle$ $\triangle$ $\times$ Secure architecture across HW/Omniscient/Application layers
[10646523] $\times$ $\times$ $\times$ $\times$ Channel modeling for communication/sensing and RCS models
[10330577] $\triangle$ $\times$ $\times$ Unified multi-modal ISAC framework
[10770127] $\times$ $\times$ $\times$ $\times$ Interference management for multi-cell ISAC systems
[10989512] $\times$ $\times$ $\times$ $\times$ RIS-enhanced physical-layer security for ISAC
[10833623] $\triangle$ $\times$ $\times$ $\times$ RIS for multi-function ISAC
This Survey Emerging issues and impacts of ISAC, including the integration of the operating spectrum across RF and optical, the evolution from single-cell to multi-cell networks, the development of multi-modal cooperation, security and privacy challenges, and progress in standardization
(*) ✓   Indicates major and systematic coverage of the corresponding dimension, typically supported by dedicated analysis or structured discussion. (*) $\triangle$   Indicates partial or incidental coverage, where the dimension is mentioned but not treated as a primary focus. (*) $\times$   Indicates that the dimension is not explicitly addressed in the survey.
In response, integrated sensing and communication (ISAC) has emerged since the mid-2010s as a unified paradigm in which wireless systems are designed to simultaneously sense the physical environment and transmit information through the co-design of shared waveforms, spectrum, and hardware platforms [9606831, 9737357]. In this context, sensing refers broadly to radar-type functions, including object detection, tracking, and recognition, as well as behavior analysis and environmental awareness. Since its formal introduction, and particularly after 2020, significant progress has been achieved in key areas such as joint signal design, multifunctional antenna systems, spectrum sharing, and edge intelligence. Indeed, ISAC is transitioning from research to large-scale deployment, driven by both industrial adoption and standardization initiatives. In the industrial domain, Huawei has advanced ISAC deployment in vehicular networks [huawei2022isac], Ericsson has extended its application to intelligent transportation and urban infrastructure [ericsson2022isac], and Qualcomm has explored millimeter-wave ISAC for autonomous driving and IIoT applications [qualcomm2022isac]. On the standardization side, 3GPP has identified 32 ISAC use cases for 6G [3gpp2022isac], ETSI’s Industry Specification Group has proposed 18 additional cases [ESTI2025isac], and IEEE finalized the first Wi-Fi sensing standard (802.11bf) in October 2024. At the global policy level, the ITU has designated ISAC as one of six core 6G application scenarios [itu2021vision], and the World Economic Forum has recognized it among the top ten emerging technologies of 2024 [wef2024emergingtech]. Collectively, these developments underscore ISAC’s accelerating momentum and its potential to become a cornerstone technology in next-generation (6G) wireless networks.
I-B Related Survey Papers
A broad collection of surveys has examined ISAC from complementary perspectives. Early contributions focused on spectrum and hardware reuse mechanisms [8999605], as well as waveform and multiple-input multiple-output (MIMO) techniques [9540344]. These works were followed by surveys on internet of things (IoT)-oriented ISAC architectures [9606831], signal-processing techniques for 5G-Advanced/6G [8828030, 10012421], channel state information (CSI)-based Wi-Fi and RF sensing [10.1145/3310194], and the historical and theoretical foundations of joint radar–communication systems [10188491, 9705498], forming the foundations of ISAC research.
Built on these foundations, subsequent surveys have expanded ISAC along several technically distinct directions. One important stream has delved into physical-layer mechanisms, including multi-carrier and dual-functional waveform construction, ISAC-tailored channel modeling, and interference-management approaches for dense or heterogeneous deployments [10770016, 9737357, 10646523, 10770127]. These studies deepen the understanding of signal construction, propagation behavior, and hardware-related tradeoffs in practical ISAC deployments. At the deployment and architectural level, reconfigurable intelligent surfaces (RIS/IRS) have been widely investigated for cooperative networking [10726912, 10702570, 10833779], with further extensions toward RIS-assisted non-orthogonal multiple access (NOMA) for enhanced joint sensing and communication (S&C) efficiency [10781415]. Application-driven surveys have addressed ISAC in vehicular networks [10418473] and uncrewed aerial vehicles (UAV)-based implementations [10004900], while the notion of perceptual mobile networks has been introduced to describe mobile infrastructures capable of dynamic S&C fusion [9296833]. Complementing analytical advances, hardware testbeds and empirical evaluations have validated ISAC concepts in practice [10049817].
Beyond physical-layer and deployment-oriented studies, another stream of literature explores ISAC’s integration with artificial intelligence (AI), security, and emerging perception paradigms. These efforts span adaptive and environment-aware transmission, secure IoT architectures, and data-centric learning frameworks for ISAC [11015430, 10989512, 10608156]. Recent studies on multi-view sensing, generalized architectural principles, and emerging methods for multi-modal S&C integration [10330577, 10049304, 10536135, 9585321] highlight the ongoing shift toward richer environmental awareness and learning-driven ISAC capabilities.
Viewed collectively, these surveys offer substantial depth across specific ISAC dimensions—waveforms, channel modeling, network cooperation, RIS-enhanced operation, multi-modal perception, security, and AI-assisted inference. However, their viewpoints remain inherently theme-centered: each provides a focused treatment on a subset of these dimensions, typically around a single dominant theme (e.g., security-centric, RIS-centric, or multi-modal-centric), but none attempts to synthesize how ISAC has evolved simultaneously across spectrum expansion, network-scale development, sensing-modality diversification, system-level optimization and security constraints, and progressing standardization.
As summarized in Table I-A, the existing surveys form a collection of rich but isolated slices of the broader ISAC landscape. A consolidated multi-dimensional evolutionary perspective that connects spectrum expansion, network-scale development, sensing-modality diversification, system-level optimization and standardization into a coherent trajectory is still not fully developed, even as such a perspective becomes increasingly important for guiding ISAC toward heterogeneous, scalable, and deployment-ready 6G systems.
Figure 1: Structure of the Survey
I-C Contributions
The main contributions of this article are as follows.
  • Sensing from RF to Optical: We present a comparison between RF-based and optical ISAC, highlighting technical challenges and synergies and emphasizing the role of advanced antenna technologies and hybrid RF-optical architectures in enhancing 6G performance.
  • Network Architecture from Single-Cell to Multi-Cell: We examine ISAC evolution from single-cell to multi-cell deployments, emphasizing waveform design innovations, inter-network collaboration strategies, and large-scale deployment, while addressing challenges such as synchronization, interference management, and resource allocation.
  • Sensing Modalities from Single-Modal to Multi-Modal (Edge Perception): We trace the progression from single-modal to multi-modal sensing, with a focus on AI-enabled real-time processing, adaptive learning, and the development of high-quality datasets essential for ISAC optimization and benchmarking.
  • ISAC Security and Privacy Enhancement: We investigate vulnerabilities in both communication and sensing domains, highlighting novel risks stemming from their dual-domain integration, and discuss opportunities for leveraging joint S&C functionalities to enhance system-level security.
  • Standardization Progress: We synthesize progress across major standardization bodies, including 3GPP, IEEE, and ITU, detailing advancements in New Radio (NR) and Wi-Fi sensing integration, spectrum management strategies, and emerging application domains such as low-altitude network operations.
Taken together, these contributions deliver a unified evolutionary framework that connects previously disjointed research themes into a coherent system-level narrative. This multi-dimensional synthesis, spanning spectrum, architecture, perception, security, and standardization, fills a key gap in existing ISAC surveys and lays the foundation for the broader adoption of ISAC within heterogeneous, scalable, and deployment-ready 6G systems.
The structure of this paper is outlined in Fig. 1. Section II compares RF-based and optical ISAC, detailing their respective challenges, synergies, and prospects for integrated RF-optical designs in 6G. Section LABEL:section3 explores the development from single-cell toward multi-cell ISAC architectures, while Section IV focuses on the evolution from single-modal to multi-modal sensing enabled by edge intelligence. Section V addresses critical security and privacy issues arising from dual-domain S&C functionalities. Lastly, Section VI reviews ongoing standardization efforts and concludes this article.
II Operating Spectrum across RF and Optical
II-A RF-based ISAC
RF technology, operating in the 3 kHz–300 GHz range, remains the foundation of wireless communication and sensing. In ISAC systems, transmitter design begins with stable carrier generation via oscillators and synthesizers, followed by modulation through mixers and upconverters to embed data or probing signals. The RF chain, comprising low-noise amplifiers (LNAs), power amplifiers (PAs), and RF filters, conditions the signal before impedance matching and directional coupling delivers it to the antenna for electromagnetic radiation. In receiver design, incident signals are captured by the antenna, amplified, downconverted, and digitized through analog-to-digital converters (ADCs) [walden2002analog], followed by digital signal processing (DSP) to recover transmitted data or extract environmental features [sturm2011waveform].
While antenna arrays, particularly massive MIMO and millimeter-wave (mmWave) configurations, enable beamforming, beam steering, and spatial multiplexing for concurrent multi-user communication and sensing, overall performance is equally constrained by RF front-end and baseband hardware. Current ISAC research emphasizes antenna architectures, suggesting a future need for deeper hardware-level performance analysis that holistically optimizes LNAs, PAs, RF filters, ADCs, and DSP modules to maximize joint S&C efficiency.
II-A1 Advanced Antenna Technologies
Leveraging spatial degrees of freedom to enhance both communication and sensing, multi-antenna technologies constitute a key enabler for next-generation ISAC. Three architectural paradigms are especially impactful: centralized arrays, distributed arrays, and movable/fluid arrays.
Centralized Antenna Architectures
These encompass compact and sparse array deployments. Compact arrays, with closely packed elements, offer high beamforming gain and spatial multiplexing capability, but at the cost of increased hardware complexity and a limited field of view. In contrast, sparse arrays employ non-uniform geometries (e.g., nested, coprime [5609222]) to synthesize large virtual apertures, reducing the number of active RF chains while retaining beamforming functionality. Although resolution and robustness depend on array design and signal processing complexity, sparse architectures strike an effective balance between cost and angular resolution, making them attractive for ISAC hardware efficiency. For instance, Zeng et al. [zeng2024fixedmovableantennatechnology] derived unified Cramér–Rao bound (CRB) and signal-to-noise ratio (SNR) scaling laws for fixed and mechanically movable uniform linear arrays (ULAs), showing that doubling the aperture halves the bearing-error variance but proportionally narrows the allowable carrier-phase mismatch. Li et al. [li2025sparse] demonstrated that uniform sparse and coprime layouts, when constrained to the same aperture as dense ULAs, preserved main-lobe width and even outperformed them in sum-rate and direction of arrival (DOA) estimation root mean square error (RMSE) at sub-0 dB SNR, using only one-quarter of the RF chains. Building on these insights, Elbir et al. [10371037] introduced deep learning-assisted hybrid beamforming that reduced the antenna selection search space by nine orders of magnitude while limiting spectral efficiency loss to under 5% across 0–20 dB SNR. Taken together, these studies underscore that sparse-geometry optimization, combined with learning-based beamforming, can achieve sub-degree angular estimation accuracy with substantially reduced hardware overhead.
TABLE II: Comparison of Antenna Technologies and Reconfigurable Metasurfaces in ISAC
()(*) $s_{0}$ is the reference waveform and $\epsilon$ controls waveform energy similarity. $F$ and $P$ represent the matrices for subcarrier allocation and power allocation, respectively. $K$ is the set of communication users. $S$ is the set of OFDM subcarriers. $\mathbf{w}$ is the receive beamforming vector. $\mathbf{H}$ represents the channel matrix. $\mathbf{R}_{\text{intf}}$ is the interference covariance matrix. $\sigma^{2}$ is the noise power. $\mathbf{I}$ is the identity matrix. $P$ is the transmit power. NN is the noise power.
Code Division
In the code division mode, S&C signals are carried by orthogonal/quasi-orthogonal sequences, occupying the same resources in the time-frequency domain to obtain higher resolution distance and speed estimation. A new code division scheme was proposed in [10202575]. The scheme used three-dimensional complementary coded scrambling to combat wireless fading and mutual interference, acting as a soft isolator for integrated radar and communication systems in multi-user and multi-target scenarios, as shown in Fig. LABEL:fig:4_divisions.
In rapidly changing environments, balancing determinism and randomness in code division is particularly crucial. Determined waveforms can improve the ability to collect environmental information for a radar system, while communication systems rely on high randomness to achieve efficient data transmission and interference resistance. Therefore, careful planning of the codeword space is necessary for designing code division systems, ensuring that radar and communication functions can complement each other and reduce mutual interference while finding a balance between sensing performance and communication efficiency.
III-A2 Unified Design
As ISAC systems evolve, the inherent limitations of orthogonal designs, such as suboptimal spectrum utilization, restricted flexibility in resource reconfiguration, and the inability to fully exploit joint processing gains, have become increasingly evident. This motivates the transition toward unified waveform design, where S&C share spectrum, hardware platforms, and processing chains under a joint resource allocation framework.
A unified waveform must balance conflicting requirements: communication waveforms are typically optimized for high spectral efficiency and data throughput, whereas sensing waveforms demand favorable ambiguity function properties and constant envelope characteristics to ensure accurate range, velocity, and angle estimation. The design problem is therefore formulated with task-specific constraints and multi-objective optimization goals, as summarized in Table LABEL:tab:waveform_constraints. Depending on the design priorities and waveform formats, three main strategies have emerged: sensing-centric, communication-centric, and joint co-design approaches.
Sensing-Centric Design
In sensing-centric designs, communication functionality is embedded into a radar-optimized waveform, with a primary focus on preserving high-fidelity sensing. A representative waveform is the chirp-based frequency-modulated continuous wave (FMCW) signal, denoted by
$s_{\text{FMCW}}(t)=\exp\left[j2\pi\left(f_{0}t+\frac{k}{2}t^{2}\right)\right],$ (1)
where $f_{0}$ is the central frequency and $k$ denotes the chirp rate. The linear frequency sweep provides fine range resolution and robustness to noise. Compared with OFDM, FMCW generally requires much lower sampling rates, which simplifies receiver processing. However, this advantage comes at the expense of significantly reduced data transmission rates, making FMCW more suitable for sensing-centric designs [sturm2011waveform].
Additionally, index modulation schemes allow information to be conveyed through antenna or frequency selection without altering the primary radar waveform. Recent studies have explored these embedding and modulation techniques, emphasizing trade-offs in symbol rate and beam fidelity. For example, [9345999] demonstrated a spatial modulation-based ISAC prototype achieving a symbol error rate (SER) below $10^{-2}$ at 15 dB SNR while maintaining angular accuracy within $1.8^{\circ}$. Similarly, [liu2020joint] proposed a joint beamforming framework for multi-user MIMO ISAC, maintaining radar gain degradation within 3 dB while achieving near-optimal communication sum rate. However, sensing-centric designs generally offer limited communication capacity and impose higher receiver processing complexity.
Communication-Centric Design
Conversely, communication-centric designs superimpose sensing functions onto existing communication waveforms, most notably OFDM and orthogonal time–frequency space (OTFS) [8288677] modulation [wei2021orthogonal, 10638525].
OFDM, widely adopted in 5G NR and IEEE 802.11 standards, provides high spectral efficiency and robustness against multipath fading by transmitting data over multiple orthogonal subcarriers. This facilitates flexible allocation of time–frequency resources for simultaneous sensing. The baseband OFDM signal can be expressed as
$\begin{split}\!\!\!\!s_{\text{OFDM}}(t)\!=\!\sum_{m=1}^{N_{s}}\sum_{n=1}^{N_{c}}d_{m,n}\,e^{j2\pi n\Delta f\,t}\,\text{rect}\!\left(\!\frac{t-(m-1)T}{T}\right),\end{split}$ (2)
where $N_{s}$ is the number of OFDM symbols, $N_{c}$ is the number of subcarriers, $d_{m,n}$ are the data symbols, $\Delta f$ is the subcarrier spacing, $T$ is the OFDM symbol period, and $\text{rect}(\cdot)$ is the rectangular function.
For radar processing, the range-Doppler matrix is constructed to achieve effective separation and joint estimation of delay and Doppler as expressed by
$Y(m,n)\!=\!\sum_{l=1}^{L}\alpha_{l}\,d_{m,n}\,e^{-j2\pi n\Delta f\,\tau_{l}}\,e^{j2\pi mT\,\nu_{l}}+W(m,n),$ (3)
where $\tau_{l}$ and $\nu_{l}$ represent the delay and Doppler shift, respectively, $\alpha_{l}$ is the reflection coefficient for the $l$-th target, $W(m,n)$ is the frequency-domain noise, and $m$ and $n$ are the symbol and subcarrier indices, respectively. Applying two-dimensional (2D) processing (e.g., a 2D FFT) to $Y(m,n)$ enables joint estimation of target range and velocity. Despite these advantages, OFDM suffers from high sidelobe levels and a poor ambiguity function, which limits its sensing accuracy. To mitigate this, [10255745] optimized subcarrier power allocation, improving sensing accuracy by 19.3% and communication capacity by 8.6%. Symbol-level precoding suppressed sidelobes by 32 dB, yielding a 21% sensing gain [10771629]. Null-space projection further achieved a 15 dB sensing gain without compromising communication quality [10661215]. Nevertheless, the sensitivity of OFDM to Doppler shifts under high-mobility conditions remains a major limitation.
To address this, OTFS has emerged as a compelling alternative for ISAC applications [10638525]. The OTFS signal can be expressed as
$\begin{split}s_{\rm OTFS}(t)=\int\int s(\tau,\nu)\cdot\phi(t-\tau)\cdot e^{j2\pi\nu t}\,d\tau\,d\nu,\end{split}$ (4)
where $s(\tau,\nu)$ is the 2D modulation symbol at time-shift $\tau$ and frequency-shift $\nu$, $\phi(t-\tau)$ denotes the pulse shape function, $e^{j2\pi\nu t}$ is the complex exponential carrier frequency modulation term. By inherently mapping information symbols to delay–Doppler coordinates, OTFS couples communication signaling with parameter estimation, resulting in a distinctive “pushpin-shaped” ambiguity function that enables high resolution in both delay and Doppler.
Recent advances have further validated the dual-functional superiority of OTFS. Multi-objective optimization has been applied to improve subcarrier adaptability and spectral reuse efficiency by 23.8% [10257601]. Waveform designs incorporating sidelobe suppression achieved up to 20 dB sidelobe reduction and a 12.4% improvement in bit error rate [10521551]. By reusing downlink echoes for uplink channel estimation under fractional Doppler, OTFS demonstrated robust performance without requiring dedicated sensing waveforms [10978674]. Delay–Doppler priors were also leveraged to refine grid resolution and reduce pilot overhead [10941992]. A unified iterative framework was later proposed to jointly perform parameter association, channel estimation, and data detection, resulting in more than 35% NMSE reduction and up to 18.2% higher detection accuracy in high-mobility environments [10409528]. These contributions collectively demonstrate that OTFS offers a promising foundation for tightly integrating S&C in dynamic wireless systems.
Joint Design
In contrast to separate optimization strategies, joint S&C design formulates the objectives of the two functions within a unified mathematical framework. This formulation enables a direct characterization of their coupled behaviors and provides a principled view of how sensing and communication interact over shared wireless resources.
The motivation for unified formulations arises from the fundamental mismatch between the waveform requirements of the two functionalities. Sensing typically benefits from deterministic and energy-concentrated signals that enhance estimation precision[he2012waveform], whereas communication favors random and spectrally efficient signaling to maximize data throughput and robustness[cover1999elements]. If optimized independently, these different requirements often yield incompatible designs, thereby motivating joint objectives that explicitly encode the shared resource constraints and physical coupling[9737357, 10418473, 9585492].
Among various unified metrics, MI is widely adopted because it simultaneously characterizes the recoverable information in communication and the observability of target-induced perturbations required for sensing. MI is also grounded in information-theoretic limits, making it resilient to practical impairments such as partial channel knowledge, interference, and clutter—conditions frequently encountered in ISAC deployments[10466681].
The joint optimization problem is typically formulated as a weighted MI maximization
$\max$ $\quad\rho I(X;Y_{c}|H_{c})+(1-\rho)I(H_{s};Y_{s}|X)\quad,$ (5)
where $X$ denotes the jointly designed transmit signal, $\rho\in[0,1]$ is a weighting factor, I denote the MI, and $Y_{c}$, $Y_{s}$, $H_{c}$, and $H_{s}$ represent the received signals and channel/target impulse responses for communication and sensing, respectively. This formulation identifies the Pareto-optimal MI frontier, illustrating the fundamental trade-offs between S&C.
Given that power and bandwidth are consumed by sensing and communication, ISAC systems inherently require a unified resource allocation framework. Sensing accuracy improves with transmit energy and bandwidth, whereas communication simultaneously requires rate, latency, and reliability guarantees. Because these demands originate from the physical coupling of a dual-functional transceiver, resource allocation emerges as a central design challenge.
$\begin{array}[]{l}\!\!\!\displaystyle\max_{\mathbf{p},\mathbf{b}}\quad\text{Sensing QoS (detection, localization, tracking)}\!\\ ~\text{s.t.}\quad R_{\mathbf{p},\mathbf{b}}\geq R,\quad\mathbf{1}_{M}^{T}\mathbf{p}=P_{\text{total}},\quad\mathbf{1}_{M}^{T}\mathbf{b}=B_{\text{total}},\end{array}$ (6)
where $R_{\mathbf{p},\mathbf{b}}$ denotes the achievable communication sum-rate, required to exceed a minimum threshold $R$. The vectors $\mathbf{p}$ and $\mathbf{b}$ represent the power and bandwidth allocated across $M$ sensing targets or communication users, subject to total resource constraints $P_{\text{total}}$ and $B_{\text{total}}$, respectively.
The sensing objectives originate from classical radar detection and estimation theory [van2002optimum], forming the fundamental limits for detecting, localizing, and tracking objects under noisy reflections. Hence, these criteria appear consistently in ISAC optimization.
  • Detection Quality of Service (QoS):
    $P_{D}=1-F_{\chi^{2}_{2}}\left(\frac{2\delta}{N_{r}\,\sigma_{r}^{2}\,\left(1+p_{q}\,\epsilon_{q}\right)}\right),$
    where $P_{D}$ denotes the probability of correctly declaring the presence of a target; $\epsilon_{q}$ is the normalized sensing channel gain associated with transmit power $p_{q}$ and array gain; $\delta$ is the detection threshold; $N_{r}$ is the number of receive antennas; and $\sigma_{r}^{2}$ is the noise variance. Physically, $P_{D}$ measures the system’s ability to detect a target under noise and clutter. A higher $P_{D}$ directly reduces miss detection in safety-critical applications (e.g., collision avoidance or intrusion monitoring), which explains why power allocation in low-SNR regimes often prioritizes reliable detection.
  • Localization QoS:
    $\text{CRB}(d_{q})=\frac{\beta_{1}}{p_{q}\,|c_{q}|^{2}\,b_{q}},\qquad\text{CRB}(\theta_{q})=\frac{\beta_{2}}{p_{q}\,|c_{q}|^{2}}$
    $\rho_{q}=\frac{\omega_{r}}{\text{CRB}(d_{q})}+\frac{\omega_{\theta}}{\text{CRB}(\theta_{q})},$
    The CRB specifies the minimum achievable variance of unbiased estimators for range $d_{q}$ and angle $\theta_{q}$. These bounds decrease with transmit power $p_{q}$, in the case of range estimation, with allocated bandwidth $b_{q}$. The constants $\beta_{1}$ and $\beta_{2}$ depend on waveform parameters and array geometry. A larger bandwidth $b_{q}$ improves delay (range) resolution, whereas a higher transmit power $p_{q}$ enhances angular resolution through improved SNR. The composite metric $\rho_{q}$ thus reflects the achievable localization accuracy of an ISAC transceiver and is essential for beam alignment, user positioning, and environment reconstruction.
  • Tracking QoS:
    $\text{PCRB}_{\xi_{q,n}}=\mathbf{J}_{\xi_{q,n}}^{-1}(p_{q,n},\,b_{q,n}),~\rho_{q,n}=\text{trace}\left(\text{PCRB}_{\xi_{q,n}}\right)$
    where $\xi_{q,n}$ is the dynamic target state (e.g., position and velocity), and $\mathbf{J}_{\xi_{q,n}}$ is the corresponding Fisher information matrix. The posterior CRB (PCRB) characterizes the tightest uncertainty bound for estimating time-varying trajectories given the sensing update rate. A smaller PCRB indicates more stable and accurate tracking, which is particularly critical in high-mobility environments. Since ISAC systems must simultaneously accommodate per-slot communication constraints (e.g., latency and throughput), the transmit power $p_{q,n}$ and bandwidth $b_{q,n}$ available for sensing directly determine the achievable tracking accuracy and update rate.
Recent studies have introduced frameworks addressing joint optimization challenges. Closed-form characterizations of ISAC metrics such as false-alarm and detection probabilities under power allocation were provided in [10124135], offering insights into how transmit energy affects sensing accuracy. A flexible resource allocation strategy for jointly distributing power and bandwidth was developed in [10387517]. Moreover, an optimal sum-of-ratios-based solution was proposed in [10217150] to coordinate power, time, and quantization budgets across distributed devices. These developments demonstrate that ISAC optimization objectives arise naturally from physical constraints, estimation limits, and shared-resource coupling, rather than being imposed by arbitrary design preferences.
Practical deployment considerations further underscore the relevance of these formulations. Hardware impairments, synchronization offsets, finite-rate control signaling, and dynamic traffic loads introduce deviations from idealized assumptions, affecting both sensing and communication. Because these non-idealities influence both functionalities simultaneously, joint S&C optimization aligns closely with the practical behavior of integrated transceivers[10418473, 10273396].
Overall, the existing studies indicate the increasing importance of developing efficient and adaptive joint resource allocation mechanisms in ISAC, which provide a foundation for scaling ISAC capabilities toward larger and more heterogeneous deployments.
Figure 4: Topologies of single-cell networks, including collaboration between BS and UEs, and BS collaboration with target monitoring terminals.
III-A3 Summary and Discussion
The initial phase of ISAC evolution is primarily characterized by the transition from orthogonal to partially integrated resource allocation, where S&C functions share spectrum resources within controlled interference levels. The feasibility of dual-functional waveforms was experimentally validated in [9903001], where real-time prototypes confirmed that practical trade-offs between the two domains could be achieved in controlled test environments. These early studies demonstrated that simultaneous transmission and sensing were technically feasible, although the coexistence of the other inevitably constrained the performance of each domain.
Subsequent research efforts shifted towards developing waveform designs capable of leveraging common hardware architectures and unified signal processing algorithms[10418473]. This stage emphasized the need for a holistic framework that not only maximizes individual link performance but also coordinates cross-domain resource utilization for global system efficiency. Nevertheless, fundamental challenges persisted, particularly in characterizing performance boundaries and in modeling the complex mechanisms of cross-domain interference. While studies such as [liu2023fundamental] made initial progress in analyzing interference in shared-spectrum ISAC systems, this area remained underdeveloped compared to traditional communication or radar theory.
From a practical standpoint, reusing communication waveforms (e.g., OFDM-based 5G NR signals) offers clear advantages in terms of compatibility and cost efficiency[8288677]. However, sensing performance remains suboptimal when relying on communication-centric designs. This mismatch underscores the necessity of developing task-driven and sensing-optimal ISAC strategies, especially under imperfect synchronization, dynamic channels, and heterogeneous task demands.
In summary, the evolution of ISAC waveform design has progressed from feasibility demonstrations to integrated optimization, establishing a foundation for practical dual-functional operation. However, key challenges remain unresolved. Interference characterization in shared-spectrum environments is still not fully understood, performance limits have yet to be rigorously quantified, and the design of sensing-optimal waveforms under realistic constraints requires further exploration. Addressing these issues is essential for moving beyond ad hoc reuse of communication signals toward systematically engineered ISAC frameworks capable of supporting diverse 6G applications.
III-B Topology Evolution from Single-Cell to Multi-Cell Network
Having established the fundamental waveform design principles within single-cell ISAC systems, our discussion was extended to multi-cell architectures, where cooperation among distributed nodes introduced additional opportunities and complexities.
III-B1 Types of Sensing Entities
BS
In ISAC networks, BSs (including macro BSs and micro BSs) provide connections between UEs and the core network. They manage wireless access, including sending and receiving S&C signals as well as controlling and managing UEs.
UE
UE are equipped with specialized sensors to gather data about the environment or surrounding objects and communicate with BSs using wireless signals. Owing to mobility and flexibility, multiple UEs collaboratively extended sensing coverage. For example, satellites in the global navigation satellite system (GNSS) and BSs in current cellular networks can utilize device-based sensing technology to locate mobile devices [liu2022networked].
Additional UE
Additional UEs correspond to those terminal devices that cannot connect to the BS directly, including smartwatches and fitness trackers. Equipped with cameras, inertial, and environmental sensors, they connected to smartphones or other UEs via Bluetooth/Wi-Fi and enhanced area awareness and mobility coverage at low cost and with high portability.
Summary and Discussion
Multiple sensing entities cooperated to rapidly fuse multi-perspective observations. The aggregated data were delivered over backhaul to processing nodes for inference, improving detection accuracy, parameter estimation, and effective coverage. Representative topologies included single-cell, multi-cell, and space–air–ground integrated network, as detailed below.
Figure 5: Topologies of multi-cell Network, illustrating: collaboration of RRUs in C-RAN, coordination among multiple BSs for joint monostatic and bistatic sensing, integration of monostatic and bistatic sensing at a single BS, and cooperation between macro and micro BSs.
III-B2 Topology of Single-Cell Network
Single-cell topology enables rapid integration of environmental information with minimal modifications to existing infrastructure. Standalone BSs provided direct services to terminals without reliance on the legacy core network. As illustrated in Fig. 4, the BS acts as the master node, while UEs and additional UEs act as subordinate nodes for cooperative sensing within the coverage area. Full-duplex operation is assumed to support simultaneous transmission and reception, with sensing integrated to optimize resource use and network performance.
Collaboration between BS and UEs
The BS and UEs collaboratively fuse signals to enhance sensing. In downlink sensing, the BS may combine uplink probes from UEs with locally received echoes, or fuse downlink echoes with UE-side reflections. Since BS and UE are spatially separated, relative delays and Doppler shifts arise, which require estimation. Triangulation-based techniques are often applied to resolve ambiguities. Additional UEs equipped with sensors can measure vital signs or motion state. They forward processed features or raw measurements to peer UEs and the BS through sidelink reference signals. This cooperation extends the sensing footprint and enables more flexible coverage.
Collaboration of BS and Target Monitoring Terminals
Target monitoring terminals are passive sensing nodes (e.g., passive radar) without active transmission. Deployed across target areas and connected to BSs over low-latency links, they expand coverage with minimal interference.
Single-cell ISAC systems face inherent challenges. Sensing accuracy and coverage are limited to the BS’s service area, especially in dense urban scenarios. The coexistence of S&C introduces intra-cell interference, such as uplink/downlink and UE-to-UE interference, degrading signal separability and efficiency. Moreover, scalability is constrained, as a single BS may become overloaded when the sensing requirement grows.
These limitations highlight the need for tailored strategies such as interference mitigation, coordinated beamforming, and asynchronous sensing support. A key issue is intra-cell interference arising from overlapping radar and communication functionalities at centralized BSs. In [han2024cellular], a monostatic ISAC configuration was investigated, where frequency-domain null-space projection reduced interference power by more than 10 dBm while preserving sensing performance. To further improve operation, [10577579] proposed a coordinated multipoint framework incorporating distributed beamforming and joint SINR optimization, which achieved communication SINR gains of up to 8 dB without degrading radar detection accuracy. Beyond waveform-level coordination, [9349171] studied uplink sensing under asynchronous transceivers and constrained antenna resources. Cross-antenna cross-correlation accurately estimates delays and Doppler shifts without strict synchronization, effectively resolving multi-target ambiguities by exploiting spatial correlation across receive antennas.
Overall, the transition from single-cell to multi-cell ISAC systems reflected a progression from localized integration toward large-scale collaborative architectures. While BSs acted as central nodes and UEs devices extended coverage and functionality, persistent interference, scalability, and coverage limitations motivated multi-cell coordination.
III-B3 Topology of Multi-Cell Network
Multi-cell ISAC architectures extend sensing coverage and enhance spatial diversity through coordination among multiple BSs (or gNBs in 5G NR terminology), as shown in Fig. 5. Overlapping service areas enable multi-perspective data fusion, thereby enhancing robustness and granularity.
Figure 6: Topologies of Space-air-ground Integrated Network
Collaboration of Multiple BSs
In cooperative ISAC networks, multiple BSs are jointly operated in both monostatic and bistatic modes. Monostatic sensing relies on locally received echoes, whereas bistatic sensing exploits echoes from adjacent BSs or UEs, thereby forming multi-static configurations that enhance angular resolution and robustness against occlusion. To address inter-cell interference in dense deployments, coordinated beamforming and interference nulling have been proposed. For instance, [10769538] developed a stochastic geometry-based framework showing that interference suppression improved radar information rate and communication spectral efficiency by more than 20% and 15%, respectively. Beyond interference management, clustering strategies play a critical role: static clustering reduces signaling overhead, while dynamic clustering adapts to user mobility. Building upon these insights, [10969840] proposed data-aided bistatic sensing, where UEs fed back received communication symbols to assist sensing, allowing BSs to jointly process local echoes and UE-side observations.
Collaboration of Macro and Micro BSs
Heterogeneous collaboration between macro and micro BSs is particularly effective in urban and indoor scenarios, where macro BSs suffer from penetration loss and coarse coverage granularity. Strategically deployed micro BSs (e.g., femtocells) provide localized perception and support seamless indoor–outdoor integration. A hybrid-access framework in [10901518] that incorporated social trust and operator collaboration improved offloading efficiency by 35% and energy efficiency by 20%. Moreover, macro BSs can serve as MEC fusion centers to aggregate sensing data from nearby micro BSs for joint processing and beamforming optimization. Coordinated multi-BS sensing, as discussed in [10273396], has been shown to significantly improve detection accuracy and extend sensing range in dynamic environments.
Collaboration of RRUs in Cloud Radio Access Network (C-RAN)
C-RAN, introduced around 2009, further advances the multi-cell paradigm by decoupling and centralizing BS functionalities. In LTE and 5G NR, a BS (gNB) can be functionally split into centralized units for higher-layer baseband processing, distributed units for lower-layer real-time tasks, and remote radio units (RRUs) for analog transmission and reception [3gpp_tr38801, TS38401]. To address full-duplex challenges, some RRUs are designated as dedicated receivers. When combined with massive MIMO and cooperative beamforming, C-RAN enables large-scale, coordinated wide-area sensing. Joint signal- and data-level fusion in C-RAN has been shown to improve both communication throughput and sensing accuracy [9737357]. In high-density networks, optimizing cluster size, user density, and spatial resource allocation to fully leverage the performance gains [10735119].
Summary and Discussion
The progression from distributed BS cooperation to heterogeneous macro–micro deployments and centralized C-RAN frameworks highlights the growing complexity of multi-cell ISAC. While these architectures offer extended coverage, spatial diversity, and improved sensing accuracy, they also introduce challenges in inter-cell synchronization, interference management, and coordination overhead. Advanced signal processing methods, such as compressed sensing and background subtraction, have been applied to recover perception parameters and mitigate multipath distortion [rahman2019framework]. A stochastic geometry model in [10735119] demonstrated that inter-cell coordination reduced sensing interference by up to 12% and improved SINR by nearly 5 dB. Moreover, joint beamforming and SINR optimization further improved worst-case radar SNR by 8% under communication constraints [10304081]. These findings underscore the need for unified cross-layer frameworks that can balance S&C demands in dynamic multi-cell environments.
III-B4 Space-Air-Ground Integrated Network
The space–air–ground integrated network has been proposed to provide seamless connectivity and joint S&C services across multiple layers. As shown in Fig. 6, this architecture supports applications such as environmental monitoring, disaster response, and remote communications in areas where conventional infrastructure is limited. By integrating satellites, aerial platforms, and terrestrial networks, it not only extends coverage but also enables large-scale, real-time data acquisition and situational awareness across diverse geographic environments[yuan2025groundskyarchitecturesapplications].
Collaboration of Space and Ground
In space–ground collaboration, satellites in low Earth orbit (LEO), medium Earth orbit (MEO), and geostationary orbit (GEO) communicate with ground networks and stations. This collaboration enhances global situational awareness and enables sensing technologies covering entire continents and oceans. LEO satellites, located at altitudes of 160–2,000 km, provide low-latency connectivity and enable ultra-dense satellite constellations, thereby facilitating real-time Earth observation and high-frequency environmental monitoring. MEO satellites balance coverage and latency, serving global navigation, IoT connectivity, and wide-area communication. Meanwhile, GEO satellites, positioned at approximately 35,786 km, provide continuous coverage over large areas, making them well-suited for high-latency-tolerant services, such as global broadcasting and long-term environmental monitoring.
Nonetheless, several challenges hinder space–ground cooperation. Propagation delay and Doppler-induced waveform instability degrade transmission quality, as addressed in [9852292] by a hybrid analog–digital transmitter architecture mitigating beam squint in massive MIMO LEO systems. Heterogeneous channel conditions and dynamic user demands further complicate resource allocation, motivating interference-aware opportunistic scheduling schemes [10506334]. Additionally, mobility and frequent changes in topology make handover management more complex. [10118845] introduced a game-based vertical handover strategy, wherein LEO satellites facilitate dynamic resource allocation, establishing a foundation for predictive mobility in SAGIN that supports ISAC.
Collaboration of Air and Ground
In the air-ground collaboration model, drones, airships, and high-altitude platforms (HAPs) serve as intermediaries between satellites and terrestrial networks, providing flexible regional coverage and rapid deployment. HAPs at 20–50 km altitude provide quasi-stationary regional coverage, while UAVs operating from 100 m to 20 km altitude offer localized and highly adaptive service. Large UAVs (around 3,000 m) are often deployed for wide-area monitoring, whereas medium-altitude UAVs (around 1,000 m) are more effective in dense urban environments. Air–ground (AG)-ISAC systems leverage aerial mobility to enhance terrestrial perception. However, they introduce synchronization, trajectory planning, and cooperation challenges. For instance, UAV trajectory optimization was shown in [10049809] to reduce sensing interruptions and improve reliability, while [10614082] emphasized deep cooperation strategies for cross-layer resource coordination between aerial and terrestrial nodes.
Collaboration of Space and Air
Space-air collaboration is facilitated between satellites and high-mobility aerial platforms, such as UAVs and HAPs, to extend connectivity and sensing in dynamic, dispersed environments. UAVs and HAPs adjust their altitudes and positions to optimize LoS with satellites, ensuring reliable, continuous data transmission. This flexibility is critical for maritime communications, remote IoT deployments, and scenarios where maintaining LoS is challenging. Moreover, this cooperation also empowers situational awareness in dynamic environments such as disasters and urban areas. Their high mobility and adaptability enable targeted data collection, complementing satellite and ground systems to fill coverage gaps, where conventional infrastructure is compromised or absent. The space-air collaboration layer enhances perception by enabling real-time, adaptive sensor positioning for optimal data acquisition in remote or rapidly changing areas.
Summary and Discussion
SAGIN represents a next-generation framework for extending sensing resolution, coverage, and reliability across multi-layered networks. Two main deployment paradigms have been explored. The first leverages existing terrestrial infrastructures below 300 meters altitude, offering cost efficiency but limited vertical coverage and spectrum contention between uplink, downlink, and sensing functions. The second relies on dedicated low-altitude networks above 300 m, where mmWave-based fusion architectures offer high-throughput communication and fine-grained sensing capabilities tailored to specific mission requirements.
Despite its promise, SAGIN introduces unique technical hurdles. Spectrum sharing across segments is complicated by rapidly varying interference conditions, which necessitate collaborative spectrum sensing approaches, such as the distributed cognitive satellite–ground framework proposed in [hu2020energy]. Latency and real-time data processing pose additional challenges in high-mobility aerial scenarios, where distributed edge computing and predictive beamforming techniques [10745905] enable low-latency and mobility-resilient service continuity. Moreover, transmitting sensitive multi-segment data raises security and privacy concerns. Blockchain-based schemes have been proposed to guarantee tamper-proof, decentralized data integrity across space, air, and ground domains.
Figure 7: Illustration of the system architecture for the edge perception
III-C Lessons Learned
The transition from single-cell to multi-cell and eventually to SAGIN-based ISAC architectures illustrates a clear trajectory from localized integration toward large-scale collaborative networks. Single-cell deployments demonstrated the feasibility of waveform co-design but suffered from limited coverage, scalability, and interference. Multi-cell architectures addressed these issues by introducing BS coordination, macro–micro cooperation, and centralized C-RAN structures, which improved spatial diversity and sensing accuracy. Extending ISAC into the space–air–ground domain further enabled global applications such as navigation, disaster response, and environmental monitoring.
While these developments highlighted the potential of ISAC, they also introduced challenges in synchronization, inter-cell interference management, and coordination overhead. To address these, advanced waveform design, coordinated beamforming, and cross-layer frameworks are required to balance S&C demands in dynamic environments.
III-C1 How Can Clock Asynchronization Be Effectively Mitigated in Distributed ISAC Networks
Clock asynchronization remains a fundamental limitation in large-scale ISAC deployments. Misaligned oscillators introduce timing offsets, carrier frequency shifts, and phase noise, thereby degrading range–Doppler estimation and symbol detection. Research has investigated mitigation through CSI ratio-based models [10678871], CRB-based performance analysis [9617146], and advanced estimators such as iterative cross-correlation [10273396], Kalman filtering [hamilton2008aces], and optimization-based offset estimation [10130780]. Practical implementations, such as the SPACE framework for joint synchronization and positioning and anchor-assisted uplink sensing [10637442], have integrated offset estimation and compensation. Despite these advances, scalable synchronization solutions for dynamic, multi-node ISAC networks remain limited, and trade-off analyses, such as those between synchronization accuracy and resource utilization, need further development.
III-C2 How Should Sensing Architectures Be Designed to Balance Latency, Fusion Accuracy, and Scalability
Efficient sensing architecture design has been recognized as a critical enabler for the evolution of ISAC networks toward multi-cell and hierarchical deployments. At the 3GPP 6G workshop in March 2025 [6gworkshop], several candidate solutions were introduced, with a three-layer sensing framework emerging as one of the most intuitive options. This framework consists of three key components: (i) local sensing nodes, which include Remote Radio Units (RRUs), roadside units, and ISAC-enabled user equipment (UEs); (ii) intermediate aggregators, such as edge servers and distributed Baseband Units (BBUs); and (iii) the sensing function implemented within the core network. In this architecture, environmental signals are captured and pre-processed at the local nodes. The intermediate aggregators then collect and refine the measurements from multiple sources. Finally, the core network enforces system-wide requirements and carries out advanced data fusion and analytics. By distributing initial processing and partial fusion to these intermediate layers—rather than transmitting raw sensor data directly to the core—this hierarchical design reduces transmission latency, alleviates computational burdens on the sensing nodes, and mitigates the delays associated with fully centralized processing.
Looking forward, the growing density of cellular infrastructures—including terrestrial networks, vertical industries, and aerial platforms—presents a remarkable opportunity to establish ubiquitous sensing as a fundamental capability of next-generation wireless systems. By leveraging shared infrastructure, reusing spectrum, and harnessing economies of scale, ISAC can offer large-scale, cost-efficient, and resilient sensing services.
Figure 8: Interactions between AI and ISAC: AI-enabled ISAC, where AI enhances ISAC through multi-modal data fusion, efficient information exchange, and environmental understanding; ISAC-facilitated AI, in which ISAC provides precise sensing data and efficient resource management, significantly improving AI-driven signal processing and network performance; AI-ISAC Co-synergy, where AI and ISAC systems mutually reinforce each other.
IV Sensing Method from Single-Modal to Multi-Modal (Edge Perception)
The transition from single-modal to multi-modal approaches is pivotal in the evolution of ISAC systems, driven by the need for high-precision, real-time decision-making in applications such as autonomous driving, smart cities, and industrial automation. Multi-modal ISAC systems integrate heterogeneous data sources to enhance both environmental perception and communication performance. Edge intelligence accelerates this shift by enabling decentralized, real-time data processing at the network edge, reducing communication overhead, and addressing privacy concerns.
This section reviews the broadening of ISAC technologies from single-modal architectures to multi-modal integration, with particular emphasis on the interplay between ISAC and AI. By leveraging AI-driven techniques, both S&C functionalities are significantly enhanced to meet the requirements of next-generation intelligent applications.
IV-A From Edge Intelligence to Edge Perception
Modern smart environments are supported by billions of edge devices (e.g., smartphones, wearable sensors, and connected vehicles) that generate massive volumes of heterogeneous data. Traditional cloud-centric approaches, which offload computation to centralized servers, often incur high latency, excessive communication overhead, and elevated privacy risks [9815195]. Edge intelligence addresses these limitations by distributing computation across edge nodes, enabling localized data analysis and decision-making[LiTWC2020, Gartner]. Federated learning (FL), in particular, offers a privacy-preserving solution by training models locally while aggregating global insights at edge servers[Cao-ComM, 10478029].
Despite its advantages, edge intelligence faces persistent challenges. The transmission of large models and gradients can overload wireless networks, and limited channel quality (e.g., fading) impairs global model convergence. Device heterogeneity further complicates federated training, while adversarial participants may exploit vulnerabilities to manipulate sensor inputs.
IV-A1 Emergence of Edge Perception
The integration of sensing into edge intelligence gives rise to edge perception, where data are processed closer to end devices for responsive and privacy-preserving analytics [10908560]. As shown in Fig. 7, an edge perception setup typically includes three co-located modules: (i) a sensing module capturing environmental data, (ii) a computing module extracting salient features, and (iii) a communication module transmitting features to remote servers. This localized decision-making reduces latency and alleviates the burden on backhaul networks.
IV-A2 Edge Perception for Sensing Enhancement
Edge perception plays a central role in enhancing sensing capabilities through multi-modal data fusion at the network edge. By employing advanced AI algorithms, subtle patterns that elude conventional methods can be captured, thereby improving detection accuracy and situational awareness. For instance, in smart homes, sensor data is monitored to assess occupant behavior and environmental factors, which in turn facilitate the automatic optimization of temperature, humidity, and lighting to enhance comfort. In industrial IoT, real-time factory monitoring is utilized to support predictive maintenance and production scheduling, while wearable devices track patient vitals to enable prompt medical interventions. Beyond indoor and industrial settings, edge perception also plays a critical role in highly dynamic outdoor environments. In dense low-altitude operations, graph neural network–based methods enable UAVs to model local state information, extract multi-modal sensing features for conflict-risk awareness, and improve the effectiveness of autonomous avoidance decisions in large-scale swarms [liMultiscaleGraphEnhanced2025, liHierarchicalConflictResolution2025].
External sensing is conducted by edge perception systems that are equipped with on-device processing capabilities. Once sensor data are generated at distributed sources, they are gathered and analyzed locally using intelligent algorithms, which enables swift decision-making while conserving network bandwidth. Internal sensing, on the other hand, involves the continuous monitoring of network conditions (such as congestion and bandwidth constraints) to support responsive scheduling and resource allocation. For example, in intelligent transportation systems, traffic signals are dynamically reconfigured based on real-time congestion indicators, and video surveillance frameworks adjust streaming quality to maintain the integrity of critical feeds under limited bandwidth conditions.
Figure 9: Illustrate the structure of the ISAC beamforming neural network in paper[10533223]
IV-A3 Edge Perception for Communication Enhancement
In addition to sensing, edge perception strengthens communication performance by fusing diverse sensing modalities into the transmission process. Localized processing reduces end-to-end latency and alleviates computational pressure on centralized infrastructure. Multi-modal fusion improves communication reliability and resource management. In vehicular networks, radar, camera, and LiDAR inputs can be combined with RF signals to enhance interference mitigation, beamforming precision, and spectrum allocation. This synergy increases throughput and reduces latency. Beyond resource optimization, reinforcement learning and cross-attention architectures dynamically allocate spectrum resources, further improving adaptability and robustness in highly dynamic environments [shang2025energyefficientintelligentisacv2x].
IV-A4 Interplay between ISAC and AI
The integration of ISAC and AI is inherently bidirectional, with each domain augmenting the other. As shown in Fig. 8, this interplay can be understood through three key perspectives.
Edge AI-Enabled ISAC
Edge computing provides rapid model adaptation and low-latency inference for ISAC, particularly on resource-constrained devices. Unlike federated learning, which emphasizes distributed training, practical deployments prioritize lightweight edge-side inference and fine-tuning of pretrained models at the edge side.
Techniques such as neural network compression, shallow architectures, and meta-learning have been adopted to support low-overhead inference. For instance, Ahmed et al. [9467318] proposed a shallow neural network for hybrid beam selection, achieving real-time operation with a 40% reduction in computational cost while maintaining near-optimal performance in mmWave MIMO systems. Similarly, Yuan et al. [9257198] introduced a meta-learned beamforming strategy capable of rapid adaptation to new propagation scenarios with limited training samples, reducing convergence time by over 60%. Transfer learning has further enabled robust generalization, where pretrained policies improved SNR by 25% and reduced reconfiguration latency by 30% compared with baseline schemes [9367008].
These findings highlight the potential of pretrained AI models, once optimized for edge constraints, to be efficiently fine-tuned or directly deployed in ISAC systems requiring responsiveness, adaptability, and low computational complexity.
ISAC-Facilitated Edge AI
The integration of S&C within ISAC frameworks provides a synchronized, multimodal data source that significantly enhances edge AI inference. Unlike traditional architectures relying on disjointed sensor modalities, ISAC systems offer a unified and temporally aligned representation of environmental dynamics by fusing radar, LiDAR, vision, and RF features.
This synergy has been validated across multiple application domains. In vehicular networks, a transformer-based predictive beamforming scheme that jointly exploited radar and CSI data improved beam selection accuracy and link throughput under mobility, achieving up to 4.2 × gains in low-speed scenarios [zhang2024transformer]. In indoor sensing, vision transformer–based models effectively processed Wi-Fi CSI data for human activity recognition (HAR). In particular, the class-attention in image transformers architecture achieved state-of-the-art classification accuracy on the University of Texas Human Activity Recognition (UT-HAR) dataset under constrained edge inference conditions [10477406]. Similarly, lightweight CSI image-based frameworks demonstrated high recognition accuracy while requiring only 2.1 MB of memory and delivering 18 frames per second throughput on Raspberry Pi-class hardware [10190332].
Beyond these applications, ISAC-enabled multimodal observation also strengthens multi-agent decision-making in highly dynamic environments. In UAV swarms, multi-agent deep reinforcement learning methods based on spatiotemporal graph reasoning and transformer-based sequence modeling enable efficient edge-node communication and coordination, significantly improving cooperative pursuit success probabilities against highly maneuverable targets under uncertain environments [yangSpatiotemporalGraphReasoning2025, yangTransformerbasedCooperativeUAV2025, yangCooperativePursuitevasionLow2025].
As pointed out, ISAC not only enhances the quality and diversity of sensing data but also enables efficient and low-latency AI inference at the edge, thereby supporting scalable deployment in latency-sensitive applications.
AI–ISAC Co-Synergy
The interaction between AI and ISAC is inherently bidirectional, with each component reinforcing the capabilities of the other. On one hand, AI algorithms can be employed to optimize waveform selection, beamforming, and sensing strategies in real time. On the other hand, ISAC systems provide rich, temporally aligned environmental feedback, enabling AI to refine inference and decision-making processes with improved accuracy and robustness.
This co-synergistic paradigm has been exemplified in [10621049], where an ISAC-guided over-the-air federated learning framework was introduced. By leveraging sensing data quality for dynamic client scheduling and power control, the framework improved convergence robustness and reduced training latency in bandwidth-limited environments[11143883]. To further enhance adaptability, [10608156] proposed a transfer learning–based edge model adaptation framework, which enabled fine-tuning of pretrained models using few-shot local samples. This design preserved model accuracy and privacy while significantly lowering communication overhead, demonstrating the potential of ISAC-assisted adaptive learning in dynamic wireless contexts.
Looking ahead, co-optimized ISAC–AI architectures are expected to incorporate continual learning, meta-learning, and online adaptation mechanisms that dynamically update AI models in response to real-time feedback. Such integration advances the resilience and efficiency of edge intelligence and further establishes ISAC as a fundamental enabler of self-evolving, autonomous wireless systems.
Challenges and Limitations
Despite the promising synergy, the deployment of AI-driven ISAC systems remains constrained by multiple technical challenges.
A primary concern lies in the requirement for precise time and frequency synchronization across distributed S&C nodes, particularly under dynamic and mobile environments. Clock and carrier frequency offsets have been shown to significantly degrade model aggregation in federated learning and reduce the accuracy of over-the-air computation. To address this issue, distributed MIMO sensing under timing and frequency asynchronization was investigated in [10130780], where a robust localization algorithm was proposed to mitigate synchronization-induced degradation.
Another critical challenge arises from the instability of AI algorithms in non-stationary wireless environments. RL methods commonly employed for optimizing beamforming and waveform design often rely on stationary assumptions that fail to generalize under rapid channel variations and user mobility. To improve robustness, a DRL-based beamforming strategy was developed in [10571018], incorporating mobility and environmental dynamics to enhance performance in vehicular scenarios.
Energy efficiency and sensing cost further represent significant barriers in resource-constrained deployments. To alleviate these constraints, a closed-loop AI-in-the-loop framework was introduced in [cai2025aiintheloopsensingcommunicationjoint], where gradient importance sampling and stochastic gradient Langevin dynamics (SGLD)-based noise modeling were applied, achieving a 52% reduction in sensing cost and a 77% saving in communication energy. In passive sensing scenarios, a deep unfolding architecture named ISAC-NET was proposed in [10474422], alternating between channel reconstruction and target detection. This architecture yielded a 25.6% improvement in sensing accuracy and a 31.4% reduction in channel estimation error compared with conventional methods.
IV-B The Goal of the Next Phase
The development of next-generation networks aims to integrate communication, sensing, and computing into unified cognitive infrastructures that augment human intelligence. In current distributed AI frameworks, edge nodes already provide essential capabilities such as sensing, autonomy, real-time responsiveness, and adaptive learning. The long-term objective is to move beyond reactive networks—primarily focused on perception and environmental understanding—toward fully autonomous cognitive systems that support decision-making, actuation, control, and continual self-optimization.
IV-B1 The Shift towards Task-Oriented ISAC
The integration of sensing, communication, and edge intelligence represents a fundamental paradigm shift, where the emphasis shifts from maximizing individual performance to accomplishing tasks efficiently and reliably. In future networks, vast amounts of multi-sensory data with diverse spatial, temporal, and stochastic characteristics will be generated at the edge. The processing of this data will directly affect edge intelligence and overall task execution. Moreover, the relative priority of S&C services will vary depending on task requirements. Consequently, data collection and transmission need to be tightly coupled with edge computing, ensuring that QoS definitions for ISAC are redefined from a task-oriented perspective. Within this framework, edge devices are expected to deliver optimized sensing performance while operating under constrained communication and computing resources.
Figure 10: The structure in paper[xu2024penetrative]
IV-B2 Enabling AI Technologies
By leveraging advanced algorithms to process and interpret high-dimensional sensory data, AI transforms conventional networks into adaptive, context-aware infrastructures. For instance, a lightweight unsupervised neural network was proposed in [10533223] to optimize beamforming in RIS-aided ISAC systems, achieving reduced computational complexity and enhanced spectral efficiency, as illustrated in Fig. 9. Representative AI-based methods for ISAC are summarized in Table VI.
TABLE VI: Summary of Deep Learning Methods of ISAC systems
Reference Signal Processing Method Application Performance
[li2021mmwave] High-pass filtering on range-Doppler data CNN Gesture recognition Accuracy >96%
[yadav2022tinyradar] Noise filtering and data normalization Posture detection Accuracy >93%
[salami2022teslarapture] Sparse point cloud processing CNN + BiLSTM + kNN Gesture recognition Accuracy >98%
[10580677] Fourier transform for spectral analysis RNN Vehicle-to-infrastructure (V2I) networks Accuracy >95%, latency reduction >20% compared to conventional methods
[Zhou2023BFPNetADL] Kalman filtering for motion state estimation Beam tracking in IoV systems Accuracy >92%, reduced beam misalignment by 30%
[grobelny2022mm] Fast fourier transform (FFT) for radar signal processing, noise filtering LSTM Hand gesture recognition for contactless interfaces (vending machines, ATMs, etc.) Accuracy >92%, reduced beam misalignment by 30%
[10571266] Pre-processing with normalization and filtering Autonomous driving, smart cities Detection probability >80%, RMSE reduction >23.5%
[9906941] Processing of angle data with noise filtering CNN+LSTM Predictive beamforming in vehicular networks Angle prediction accuracy >15%, robust to estimation errors
[signals3020017] Cell average constant false-alarm rate algorithm, data filtering Human activity recognition (boxing, jumping, squatting, etc.) Recognition accuracy >97%, 1% higher than using single data types
[hayashi2021radarnet] Miniature radar signal processing Gesture recognition Inference time <0.15 ms, Accuracy >90%
[zhang2024transformer] Echo signal processing, spatial filtering Transformer Predictive beamforming for vehicular ISAC systems Communication sum rate >3.7 dB compared to baseline, improved performance in high mobility
[10477406] Signal denoising and phase calibration Human activity recognition (Wi-Fi sensing) Accuracy >92%
[10577431] Position-based data normalization, LiDAR downsampling Beam prediction in mmWave V2I ISAC systems Distance-based accuracy >91%, improved zero-shot robustness
[wang2022heartprint] Heartbeat-based radar sensing Supervised Learning Authentication (Heart rate data) Accuracy >97%
[10533223] Channel correlation and sensing gain maximization Unsupervised Learning Beamforming design in RIS-aided ISAC Sensing SNR >2 dB, computational complexity reduced by 50% compared to traditional methods
[10113889] Local CSI estimation, power allocation Interference management in multi-cell ISAC systems Interference mitigation >30%, reduced overhead by 20% compared to centralized approaches
[10621049] Over-the-air FL aggregation,beamforming Federated Learning ISAC networks integrating over-the-air FL for model training Reduces FL aggregation latency by 50%, mitigates sensing interference by 22%
[10669033] GAN-based data augmentation, deep Q-Network Optimization Large-Scale Multi-agent ISAC Reduces communication cost by 35%, improves FL convergence by 1.6x
[10620899] UAV beamforming, RIS phase control Multi-agent Learning UAV RIS-assisted ISAC optimization Improves sensing rate by 18%, communication rate by 22% over non-learning methods
[10571018] Vehicular beamforming and power control ISAC for Vehicle-to-Infrastructure networks Improves sum-rate by 29%, enhances sensing accuracy by 31%
Among recent breakthroughs, large language models (LLMs) have demonstrated strong potential in extracting knowledge from sensory data with minimal task-specific supervision. Typically structured into three modules, including perception, grounding, and alignment, LLMs enable multimodal reasoning and semantic abstraction[10638143]. For example, [wojcik2023beyond]introduced a framework where smartphone accelerometer signals were transformed into textual sequences for motion and environment inference, while electrocardiogram (ECG) signals were converted into text-based indicators such as heart rate, as shown in Fig. 10. Similarly, [xu2024penetrative] extended LLMs beyond textual processing to encode domain knowledge into cyber-physical systems. This signal-to-text paradigm highlights the semantic abstraction capabilities of LLMs in perception-centric ISAC networks.
However, deploying LLMs within ISAC remains a challenging task due to the latency requirements, model complexity, and hardware limitations. As noted in [11026879], cloud-based deployment supports high-level semantic reasoning but suffers from excessive inference latency (typically > 100 ms), rendering it unsuitable for real-time tasks. Conversely, edge-based deployment reduces latency (< 50 ms) and enhances privacy but requires model quantization and specialized AI accelerators (e.g., NPUs). Hybrid edge–cloud strategies provide a compromise, performing prompt-based summarization at the edge while leveraging cloud resources for advanced reasoning. Mapping deployment strategies to ISAC use cases reveals that real-time tasks such as beam prediction and mobility management benefit from edge or hybrid solutions, while latency-tolerant tasks like semantic reasoning or long-term optimization prefer cloud-based solutions. KPIs include inference delay, power consumption, model size, and Top-$\kappa$ accuracy in sensing-driven tasks (e.g., ChatGPT-4 achieving 78.4% Top-1 and 94.1% Top-3 accuracy in beam prediction [10829757]).
Although challenges regarding interpretability, energy efficiency, and real-time constraints remain, LLM-enabled ISAC frameworks offer a promising pathway toward scalable and intelligent S&C infrastructures in next-generation wireless networks.
IV-C ISAC Datasets
The design of communication modules is inherently coupled with wireless channel characteristics, which are typically analyzed using large-scale channel measurements. Similarly, the design of sensing systems relies on multimodal data collected from diverse sensors. Therefore, constructing comprehensive sensing datasets is of critical importance for advancing edge intelligence research and enabling the effective application of AI in ISAC.
Inspired by ImageNet [5206848], recent efforts have proposed systematic methodologies for building large-scale sensing datasets. The process generally involves four key steps: 1) Defining dataset objectives to guide categorization and ensure diversity; 2) Collecting massive multimodal data from heterogeneous sources such as LiDAR, cameras, mobile devices, and wireless sensors; 3) Cleaning and labeling data through automated or crowdsourcing approaches with confidence thresholds and duplicate removal; 4) Making datasets publicly available via cloud platforms to foster community contributions and benchmarking.
IV-C1 High-Quality Datasets
Several representative datasets have been developed to support ISAC and edge intelligence research:
  • SDP[sdp8]: Launched in February 2024 by IEEE ISAC-ETI, the SDP platform is the industry’s first large-scale field-measured sensing dataset. This platform addresses major wireless sensing challenges by publicly releasing RF datasets and promoting high-performance and reliable wireless sensing solutions. The platform has released a large-scale field-measured dataset that covers 7 typical sensing application scenarios. It spans 400 hours, contains 14.1T of data, and can be used for algorithm research and model training for various sensing tasks. These tasks include detection (presence detection of targets like people and pets), positioning & tracking (tracking of people, vehicles, and drones), recognition (gesture and action recognition), vital signs (breathing, heartbeat, sleep), and imaging (human skeleton and environmental reconstruction).
  • ImgFi[10190332]: The dataset was collected in indoor environments, covering diverse activities. The approach transforms CSI amplitude and phase into image-like representations, enabling deep learning-based HAR models with reduced computational overhead. The framework aims to enhance HAR accuracy while maintaining lightweight deployment.
  • WiMANS[huang2025wimans]: A Wi-Fi-based multi-user activity sensing benchmark, WiMANS provides 9.4 hours of dual-band CSI synchronized with RGB videos of up to five users performing nine daily activities. It targets multi-user activity recognition, tracking, and pose estimation, addressing the limitations of single-user datasets.
  • Radar signatures of human activities[UnivGlasgow]: This dataset, provided by the University of Glasgow, focuses on radar-based human activity recognition. It includes a range of human activity radar signals collected using FMCW radar, supporting the development and evaluation of radar signal processing and machine learning algorithms for human activity classification.
  • EyeFi[9183685]: EyeFi fuses CSI with camera data for human sensing, motion tracking, and identification tasks. It uses a three-antenna Wi-Fi chipset to estimate the Angle of Arrival (AoA) of Wi-Fi signals combined with overhead camera trajectories. Synchronization aligns CSI-based and vision-based trajectories. EyeFi performs trajectory matching to identify individuals by linking their Wi-Fi MAC addresses and physical movements.
  • OPERAnet[bocus2022operanet]: OPERAnet is a multi-modal dataset designed for indoor Human Activity Recognition (HAR) and localization. The dataset contains approximately 8 hours of data collected across two indoor environments with six participants performing six daily activities. It is designed to facilitate multi-modal research in smart homes, elderly care, and surveillance, with high-resolution activity and location labels.
  • MmWave Gesture Dataset[liu2021m]: MmWave-gesture-dataset features over 56,420 tracking samples and 1,357 minutes of gesture data, collected from two scenarios with varying distance ranges. It includes both preset and unexpected gestures, offering multi-dimensional data from raw signals to advanced processing information, which aids in gesture recognition technology development.
    TABLE VII: Summary of High-quality Datasets in ISAC
    Dataset Name Data Types Application Scenarios Data Volume Provider/Founder
    Detection Positioning Tracking Recognition Imaging
    SDP[sdp8] • CSI
    • Kinect data Size >14.1 TB Time >400 hours IEEE ISAC-ETI
    ImgFi[10190332] • CSI images $\times$ $\times$ $\times$ $\times$ Size >10,000 Nanjing Forestry University
    WiMANS[huang2025wimans] • CSI
    • RGB video $\times$ $\times$ $\times$ $\times$ Time >9.4 hours Imperial College London
    Radar Signatures of Human Activities[UnivGlasgow] • Radar $\times$ $\times$ $\times$ $\times$ - University of Glasgow
    EyeFi[9183685] • CSI
    • RGB video $\times$ $\times$ $\times$ - University of North Carolina
    OPERAnet[bocus2022operanet] • CSI
    • Kinect data $\times$ $\times$ $\times$ $\times$ Size >8 hours Bocus et al.
    MmWave Gesture Dataset[liu2021m] • mmWave
    • Camera
    • Depth maps $\times$ $\times$ $\times$ $\times$ Size >56,420 Time >1357 minutes Beijing university of Post and Telecommunications
    DeepSense 6G [deepsense6g] • RGB
    • CSI
    • LiDAR
    • GPS
    • Radar
    • mmWave $\times$ $\times$ $\times$ Size >1 million data points Arizona State University
    WALDO[nist-waldo] • CSI
    • LiDAR
    • RGB
    • Motion sensors $\times$ $\times$ $\times$ $\times$ - National Institute of Standards and Technology
    M3SC datasets [10330577] • LiDAR
    • RGB
    • Depth maps
    • mmWave $\times$ $\times$ $\times$ Size >1500 scenarios School of Electronics, Peking University
  • DeepSense 6G[deepsense6g]: Designed to support research in telecommunications and sensory technologies for 6G networks, DeepSense 6G dataset provides a rich collection of real-world, multi-modal data. It captures communication and sensory inputs across diverse scenarios and times of day, facilitating deep insights into 6G technology performance and adaptability.
  • WALDO[nist-waldo]: Representing a significant advancement in indoor positioning research, the WALDO dataset combines IEEE 802.11ay PHY software with multi-static radar systems. It provides radar and wireless signal data for developing advanced positioning algorithms capable of accurately estimating human positions in complex indoor environments.
  • M3SC dataset[10330577]: The M3SC dataset(Mixed Multi-modal Dataset) integrates data from cameras, LiDAR, GPS, and environmental sensors to provide a comprehensive view of different scenarios, including urban and rural environments and varying weather conditions. It supports tasks such as positioning, tracking, environmental monitoring, and cross-modal analysis, aiding in the optimization of system performance and the development of algorithms for adaptive resource management and energy efficiency.
    IV-C2 Summary and Discussion
    Simulation datasets provide controllable environments for analyzing system performance and optimizing algorithm design. However, they often fail to capture the complexity and dynamics of real-world environments. By contrast, field-measured datasets such as SDP and DeepSense 6G, though resource-intensive to construct, provide higher fidelity, diversity, and representativeness. Such datasets are indispensable for training AI models that generalize well to practical ISAC scenarios, thereby accelerating progress in ISAC and edge intelligence technologies[wang2025surveywifisensinggeneralizability].
    IV-D Lessons Learned
    Despite notable application-driven advances, substantial theoretical and practical challenges persist in unifying edge intelligence with ISAC. This raises several fundamental questions:
    IV-D1 How to Build Hardware That Unifies Sensing, Communication, and Computing
    Current BS architectures are inherently complex, and the direct integration of sensing functionalities is neither cost-efficient nor scalable. Sensing modules often demand high-resolution and high-precision devices, which substantially increase implementation cost and design complexity. In parallel, AI algorithms typically rely on specialized accelerators such as GPUs and NPUs, further complicating their integration into low-power, resource-constrained communication hardware. As a result, achieving hardware platforms that seamlessly harmonize communication, sensing, and computing functionalities remains a critical frontier for future research.
    IV-D2 How Can Massive, Heterogeneous Sensing Data Be Integrated at Scale
    The fusion and processing of heterogeneous, multi-dimensional sensing data demand new algorithmic frameworks that can adapt to device diversity and coupled application requirements. Within a single coverage area, sensing data are inherently interdependent, necessitating efficient mechanisms for distributed model parameter generation. However, current distributed learning approaches often overlook device heterogeneity, resulting in inefficiencies and degraded performance. In massive machine-type communication (mMTC) scenarios, resource-constrained sensors may lack the capacity to perform local training, which forces trade-offs between S&C under stringent resource budgets [ullah2018information]. Developing scalable, adaptive, and resource-aware data fusion strategies remains an open and critical challenge.
    Figure 11: Communication Security and Sensing Privacy in ISAC Systems
    V ISAC Security and Privacy
    To fully explore the potential of ISAC technology in next-generation networks, the data security and sensing privacy should be properly addressed, as illustrated in Fig. 11. The integration of S&C functions introduces new vulnerabilities due to dual-domain exposure. While sensing enhances communication, it may also inadvertently exposes sensitive information, such as the location and movement patterns of targets. Furthermore, the shared infrastructure supporting both S&C increases the risk of data interception, thereby heightening the likelihood of security breaches. Nonetheless, this integration also opens opportunities for designing enhanced security and privacy mechanisms[yang2025privacy].
    V-A Communication Security
    Integrating communication data into probing ISAC signals introduces vulnerabilities, especially when eavesdroppers intercept these signals. The risk increases when high-power signals are used for target sensing. At the MAC layer, the shared data pathways between S&C raise the possibility of routing integrity issues. Attackers may exploit sensing data for spoofing or impersonation, while the dual-use nature of ISAC amplifies the risk of man-in-the-middle attacks. At the physical layer, dual-functional beams used for both S&C make communication data vulnerable to malicious eavesdropping, as targets in the beam’s path can intercept embedded communication signals.
    Physical layer security (PLS) remains vital in ISAC systems, even if communication data is encrypted at higher layers. PLS offers additional protection against low-level threats and strengthens overall system security by preventing attacks that could compromise communication channel data. Traditional secure beamforming approaches become insufficient because they may compromise sensing functionality. This necessitates the development of novel, tailored security approaches.
    Several countermeasures to address communication data security in ISAC have been proposed.
    V-A1 Artificial Noise (AN) Design
    AN enhanced PLS by injecting interference to degrade the SINR at eavesdroppers, thus reducing their decoding ability while preserving communication and sensing quality for legitimate users. In [9199556], an optimization framework for AN-aided secure transmit beamforming in multi-user MISO ISAC systems maximized the secrecy rate by jointly designing beamformers and AN. It characterized the trade-offs between secrecy rate, communication quality, and sensing functionality and demonstrated robustness under imperfect S&C CSI, confirming the effectiveness of AN in mitigating communication security threats.
    V-A2 Directional Modulation (DM)
    DM uses symbol-level precoding to apply constructive interference (CI) to legitimate receivers while limiting signals at potential eavesdroppers by employing destructive interference (DI). [9737364] applied the CI-DI approach to mmWave frequencies, where angular correlations between S&C channels are critical. This technique effectively enhances PLS by restricting signal leakage toward unauthorized directions, reducing energy consumption and system cost in high-frequency ISAC deployments.
    V-B Sensing Privacy
    The integration of sensing functionality into wireless systems introduces additional privacy risks. Sensing signals emitted by dual-function BSs are reflected by targets. It may be intercepted by unauthorized receivers and exposing sensitive target-related information and compromising privacy. Unlike communication signals, which benefit from higher-layer encryption, passive sensing signals cannot be effectively encrypted, presenting a unique security challenge. Moreover, measures to mitigate interference between S&C functions can inadvertently reveal critical transmitter information, such as location. Enhanced sensing capabilities further elevate the risk of passive eavesdropping by unauthorized nodes. Additional concerns arise in imaging and environmental sensing, where unencrypted control signals (e.g., pilot signal) may be exploited to infer user positions or construct environmental maps. These vulnerabilities underscore the need for robust privacy-preserving mechanisms.
    Several solutions have been proposed to mitigate sensing privacy risks.
    V-B1 Beamforming for Privacy
    Innovative beamforming techniques are crucial for preventing unauthorized data extraction while maintaining S&C functionality. In [10587082], a beamforming-based method was proposed to secure sensing functionality by integrating AN. It optimized the beamforming and AN covariance matrix to enhance MI for legitimate radar users, while deteriorating it for unauthorized ones. This successfully restricts the adversary’s ability to intercept sensing data, protecting environmental privacy without degrading sensing accuracy. Furthermore, [10574259] proposed an AI-assisted adaptive beamforming method to obfuscate signals, blocking unauthorized sensing while keeping legitimate activities functional. This method suppresses unauthorized sensing while maintaining the effectiveness of legitimate sensing operations.
    V-B2 CSI-Based User Location Anonymization
    This method focuses on mitigating unauthorized positioning or tracking by obfuscating CSI changes. The goal is to obscure the user’s location while preserving sensor accuracy. In [Wang2024GenerativeAB], an AI-based signal generation method is proposed to prevent unauthorized devices from extracting CSI-based location data, resulting in a 70% reduction in the risk of inference. Similarly, [10556606] proposed a multi-antenna signal-masking technique that exploits CSI time correlation and controls phase differences to thwart eavesdropping. Additionally, [10611746] introduced a pilot-shielding method with tailored spectral characteristics to conceal genuine CSI variations, misleading eavesdroppers, and safeguarding location privacy.
    V-C Sensing-Assisted Communication Security
    While the integration of sensing functionality in ISAC systems introduces new challenges in ensuring communication security and protecting sensor privacy, it also provides unique opportunities to enhance PLS. Leveraging its sensing capabilities, ISAC systems can dynamically estimate the sensing channel, i.e., the wiretap channel, by determining the angle of arrival of each target. This enables the system to prevent the target from decoding confidential information, thereby overcoming the traditional difficulty of obtaining eavesdropping CSI, a well-established limitation in conventional communication-only systems.
    V-C1 Sensing-Assisted PLS
    Sensing-assisted PLS represents a promising approach for enhancing the security of ISAC systems, where the sensing functionality aids in improving the system’s ability to thwart eavesdropping and information leakage. By dynamically estimating the wiretap channel, ISAC systems can precisely track eavesdroppers’ locations and adapt communication strategies accordingly. This is particularly beneficial in overcoming the challenges posed by traditional methods that require difficult-to-obtain eavesdropping CSI. This allows for real-time estimation of eavesdropper angles, enabling secure beamforming strategies that degrade the signal at the eavesdropper’s location while boosting it for legitimate users. Such methods, as proposed in [10349846] and [xu2024sensingaidednearfieldsecurecommunications], explore various optimization frameworks that balance communication performance with security objectives, demonstrating the mutual benefits of S&C functionalities in enhancing PLS. In [9838753], the iterative optimization of beamforming and AN strategies results in improved secrecy rates by minimizing the potential for eavesdropping while dynamically adjusting to the eavesdropper’s estimated location. By combining S&C capabilities, these systems represent a significant advancement in securing the wireless transmission of sensitive information in next-generation ISAC networks.
    V-C2 Sensing-Assisted Covert Communication
    Unlike defending information eavesdropping in PLS, covert communication aims to prevent adversaries from detecting the existence of communication itself, requiring high-level confidentiality protection. Sensing-assisted covert communication leverages the dual functionality of radar and communication systems to ensure that information transmission remains undetectable by adversaries. These systems use sensing to track eavesdroppers and prevent communication signal detection by optimizing transmission characteristics, such as beamforming and power allocation. One approach involves jammer-assisted techniques, where a jammer is deployed to obscure transmission while the radar detects targets and communicates covertly. Optimization methods such as SDR [10507294] maximize covert throughput while maintaining radar detection constraints. Joint radar and communication beamforming designs also optimize covert throughput while maintaining radar performance. However, using dual-function waveforms introduces trade-offs between radar detection and covert transmission, balancing communication secrecy with efficient target detection. These approaches highlight ISAC’s potential to offer secure, covert communication by balancing the competing demands of S&C performance.
    V-D Lessons Learned
    Incorporating sensing data into security frameworks creates significant opportunities for co-design. However, several unresolved issues remain that warrant further investigation.
    V-D1 How to Model and Defend Against Hybrid Cross-Domain Attacks
    While extensive studies have investigated security and privacy in ISAC systems, few have systematically addressed hybrid threats, where adversaries simultaneously exploit vulnerabilities in both S&C domains. Research on sensing-assisted eavesdropping [10757770] highlighted this challenge, yet comprehensive attacker models that jointly capture sensing deception and communication interception remain scarce. Unlike conventional wireless security frameworks, ISAC requires hybrid security models that explicitly account for adversaries capable of manipulating both physical- and digital-layer information in a coordinated manner.
    V-D2 How to Validate ISAC Security Frameworks under RealWorld Constraints
    Most existing studies have relied on mathematically intensive optimization strategies, typically assuming perfect knowledge of attacker locations, CSI, and environmental conditions. In reality, ISAC systems must operate with imperfect CSI, mobility, hardware impairments, latency constraints, and limited sensing precision. Although [10529955] investigated UAV-assisted ISAC security, its experimental validation remains limited, highlighting the need for prototypes, datasets, hardware testbeds, and standardized evaluation procedures.
    Figure 12: Brief Review of 3GPP Standards Progress in ISAC
    VI Standardization Progress from Third Generation Partnership Project (3GPP) to ITU
    Standardization has played a pivotal role in shaping the ISAC landscape, addressing interoperability challenges, enabling widespread adoption, and accelerating technological innovation.This section reviews recent progress in ISAC standardization, with particular focus on the collaborative roles of 3GPP, IEEE, and ITU. Key contributions on NR air interface advancements, emerging technologies, and regulatory frameworks are also examined.
    VI-A Efforts and Trends in Standardization
    The IEEE, alongside key organizations such as the ITU and 3GPP, has played complementary roles in the development and standardization of NR air interface technologies.
    VI-A1 Brief Review of 3GPP Standards
    3GPP has progressively incorporated ISAC-enabling functionalities into its release-based standardization framework, thereby establishing the technical foundation for converged S&C systems, as illustrated in Fig. 12.
    Release 15: Foundational NR Technologies
    Rel-15 marked the launch of the first complete standalone NR architecture. Foundational technologies, including beamforming, massive MIMO, and flexible spectrum utilization, were introduced. These advancements, outlined in standards such as TS 38.300 and TS 38.104 [TS38300, TS38104], laid the groundwork for ISAC by enabling efficient resource allocation and high-performance signal processing. Although ISAC was not explicitly addressed, these technologies provided critical building blocks for future integration.
    Release 16: Positioning and Multicast for ISAC
    Rel-16 brought improved system efficiency, expanded use cases, and additional spectrum for 5G. Though it did not directly address ISAC, it introduced key enablers. Notably, 3GPP TR 38.855 introduced advanced positioning methods, including multi-round trip time (Multi-RTT), uplink-AoA, and downlink angle-of-departure (DL-AoD))[TR38855], improving accuracy in both indoor and outdoor environments. The positioning reference signal (PRS) and sounding reference signal (SRS) refined time-of-arrival and uplink-based positioning, with Fig. 13 showing the Multi-RTT procedure. Rel-16 also added multicast capabilities for vehicular use, enabling C-V2X communication. Initial mmWave support in Rel-16 paved the way for ISAC’s evolution in future 5G and 6G systems.
    Release 17: mmWave ISAC Enhancements
    Rel-17 introduced advanced MIMO, improved coverage, spectrum extension, and satellite integration, forming the foundation for comprehensive 5G upgrades. In vertical industries like factory automation and transportation, 3GPP TS 22.104 imposed stringent positioning requirements, including 20–30 cm accuracy[TS22104]. Rel-17’s 5G mmWave systems achieved 400-MHz bandwidth, offering 0.1875-meter resolution[shi2022device]. Multi-antenna BSs exploit AoA/AoD estimation technologies such as multiple signal classification (MUSIC) for precise target detection. ISAC takes advantage of the extensive cellular infrastructure and high-frequency bands, reinforced by ultra-large-scale MIMO for increased bandwidth and superior angular resolution. By combining low-latency, high-precision communication with advanced sensing, ISAC facilitates remote object detection and environmental analysis without physical contact.
    Release 18: 5G Advanced and Intelligent Positioning
    Rel-18 introduced 5G Advanced, featuring RAN intelligence and carrier-phase–based positioning, which significantly improved sensing accuracy. Key enhancements included sidelink positioning, enabling vehicles and devices to collaborate in localization in complex environments. The 3GPP TS 38.355 specification defined mechanisms for exchanging positioning information between UEs and networks, thereby enhancing V2X localization performance [3gpp-ts-38.355, 10517597]. This release also addressed multi-device coordination and dynamic resource allocation, laying the groundwork for more advanced ISAC integration in subsequent releases.
    Release 19 and Beyond: Toward ISAC Standardization
    In early 2022, Rel-19 SA1 initiated a feasibility study on ISAC services, identifying use cases and requirements for commercial ISAC implementation across industries. By June 2023, the 3GPP SA plenary organized a workshop for Rel-19 SA2, where ISAC was proposed by several companies and included as a potential study item for Rel-19. TR 22.837 [3GPP-rel19] identified 32 ISAC use cases across five deployment scenarios: smart transportation, smart industry, smart home, smart city, and smart drones.
    In August 2023, 3GPP introduced TS 22.137, Service Requirements for ISAC, during the SA1#103 meeting [TS22137]. This specification defined eight KPIs for wireless sensing in 5G, including positioning accuracy, velocity estimation, and sensing resolution. Subsequently, in December 2023, the RAN#102 meeting approved a dedicated study item on ISAC channel modeling, which was further refined at RAN#103 in March 2024 [R1-240]. The objective of this study is to establish a unified framework for object detection and tracking, thereby improving the reliability of ISAC across diverse scenarios such as UAVs, pedestrians, vehicles, and hazardous objects. Building upon these standardization efforts, ISAC air interface technologies are expected to be systematically investigated in Rel-20 beginning in 2025, with final specifications anticipated in Rel-21 as part of the 6G standardization cycle [lin2023bridge6g5gadvancedevolution].
    Summary and Discussion
    3GPP has made substantial progress in advancing converged S&C technologies for advanced 5G systems. As NR technology continues to evolve, leveraging existing cellular infrastructure for perception functions offers clear benefits, including reduced deployment costs, accelerated rollout, and efficient spectrum reuse.
    Figure 13: Multi-RTT positioning procedure: the messaging between the LMF, the gNBs, and the UE to perform the LMF-initiated location information transfer procedure for Multi-RTT.
    Enhancements in 5G infrastructure and air interface designs are expected to significantly strengthen wireless channels, thereby supporting more accurate and fine-grained environmental sensing. A representative example is Huawei’s proposal to integrate the sensing function into the 5G-Advanced core network, which aims to enable functions such as sensing control, AI-based data processing, secure data handling, and monetization of sensing data [Huawei2023]. At the same time, the design of ISAC service processes and protocol stacks remains an open challenge. Interoperability requires well-defined signaling procedures for sensing capability negotiation, demand triggering, and result reporting. Whether existing protocol stacks can be extended to accommodate these new requirements, or if entirely new signaling frameworks must be developed, remains an open research problem. Addressing these issues will be essential to ensure seamless integration of ISAC into commercial deployments.
    VI-A2 Brief Review of IEEE Standards
    In parallel with 3GPP, IEEE has made significant strides in ISAC standardization. A major milestone is the IEEE 802.11bf standard, which incorporates sensing functionalities into the existing Wi-Fi ecosystem and contributes to the broader ISAC framework. Managed by the IEEE 802.11 Task Group bf, this effort aims to formalize Wi-Fi-based sensing within the IEEE 802.11 family, complementing 3GPP’s efforts [kaushik2024toward].
    Problem Identification and Framework Development
    In 2018, IEEE began exploring Wi-Fi sensing by leveraging CSI and the received signal strength indicator (RSSI) for environmental detection. Early studies demonstrated the feasibility of using reflected Wi-Fi signals to detect motion, presence, and physiological parameters without the need for additional hardware. Key use cases included indoor localization, smart home technologies, healthcare surveillance, and gesture recognition. The IEEE 802.11bf was instituted in 2020 to standardize Wi-Fi sensing. The initiative aimed to integrate sensing capabilities into existing Wi-Fi standards while ensuring backward compatibility and minimizing resource overhead so that communication performance is preserved. It also improved the ambiguity function through optimized waveform design and reflection analysis, leading to enhanced sensing accuracy.
    Draft Development and Technical Enhancements
    By 2022, the IEEE 802.11bf draft standard had begun defining several foundational elements for Wi-Fi sensing. It formalized the use of existing Wi-Fi waveform components, including preamble training fields and pilot tones, to enable sensing without modifying the underlying frame structure, and it introduced sensing measurement reports to support the exchange of channel information for environmental inference. Refinements to OFDM-based waveforms improved sensing accuracy while preserving throughput, and enhancements to the ambiguity function enabled more reliable multi-target separation. The draft proposed mechanisms for interference mitigation, outlining dynamic strategies for allocating spectrum and time resources so that sensing operations and data transmission could coexist efficiently.
    A comprehensive framework was defined, introducing dynamic resource allocation and optimized waveform designs to support dual S&C functionalities.
    Refinement and ISAC Integration
    The IEEE 802.11bf is refining the draft standards, with final approval anticipated in 2024. Efforts are concentrated on enhancing channel measurement techniques for precise location-based services and on aligning with adjacent fields such as autonomous systems and extended reality. In addition, multi-user resource management is being optimized to support dense deployment scenarios, which are critical for robust and scalable ISAC applications.
    Summary and Discussion
    The IEEE 802.11bf standard has substantially advanced ISAC standardization by exploiting existing Wi-Fi features for dual-use S&C. Dynamic resource allocation mitigated interference challenges, ensuring reliable performance for both functions. Moreover, multi-device collaboration frameworks extend sensing coverage and accuracy, laying the foundation for scalable ISAC solutions in future Wi-Fi and beyond systems.
    VI-A3 Brief Review of ITU Standards
    In parallel with the efforts of 3GPP and IEEE, ITU has primarily focused on spectrum management and regulatory alignment, with particular emphasis on spectrum allocation and the development of frameworks that facilitate ISAC deployment in higher-frequency bands.
    Spectrum Allocation and Management
    At the World Radiocommunication Conference 2019 (WRC-19), ITU allocated higher frequency bands, including mmWave spectrum in the 24–52.6 GHz range, to support 5G. These allocations are highly favorable for ISAC, as the wide bandwidth and short wavelengths enable fine sensing resolution. In the next phase, ITU has also initiated studies on sub-terahertz (sub-THz) and terahertz (THz) bands (100-300 GHz), which are expected to be critical for high-resolution S&C in 6G networks [kaushik2024toward].
    IMT Framework and ISAC Integration
    ISAC has been formally incorporated into the IMT-2030 framework as part of ITU’s 6G vision. This framework focuses on dynamic spectrum sharing and radar–communication coexistence, enabling advanced ISAC deployments.
    Summary and Discussion
    ITU’s contributions center on spectrum allocation, global harmonization, and aligning standards with ISAC objectives. Prioritizing smart spectrum coexistence and dynamic sharing, ITU advanced frameworks support ISAC integration. ITU has also initiated discussions on the role of metasurfaces and RIS, highlighting their potential to enhance sensing resolution and optimize the performance of ISAC systems.
    VI-B Lessons Learned
    Standardization has advanced ISAC from a conceptual vision to a pre-commercial reality, yet several technical and organizational challenges remain. Insights from current standardization efforts, aligned with emerging research demands, point to the following key questions:
    VI-B1 How Can ISAC Performance be Validated through Standardized KPIs
    Standardization bodies have defined performance metrics and testing methodologies to enable systematic evaluation of ISAC. IEEE 802.11bf specifies sensing KPIs such as velocity (0.1–0.4 m/s) and angular resolution (1°–8°), which are critical for assessing the trade-offs between S&C performance. 3GPP further incorporates these KPIs into channel models to validate ISAC under real-world cellular environments, particularly in vehicular networks and IoT systems [10251772]. These benchmarks ensure that standardized ISAC solutions achieve acceptable trade-offs between sensing accuracy and communication reliability.
    Figure 14: ISAC-based rotary-wing UAV Scenario [10529184]
    VI-B2 How Can Use Case-Driven Research Guide Standardization
    3GPP SA1 has consolidated extensive research on ISAC scenarios within the RAN1 channel-modeling project, with industry consensus highlighting the pivotal roles of the low-altitude economy, drones, and intelligent transportation. For example, enhancements in 5G NR Rel-18 enable drones to support beamforming, allowing BSs to detect unauthorized flights, perform obstacle avoidance, and guide trajectories via multi-station collaboration. Integrating passive tracking with active sensing improves positioning accuracy, while environmental awareness assists beam management and mobility. Furthermore, mmWave operation achieves low latency with slot lengths as short as 0.125 ms, ensuring real-time responsiveness.
    The low-altitude economy has also gained strong policy support. In China, the Civil Aviation Administration has promoted 5G NR–based applications, and the Ministry of Industry and Information Technology announced on October 23 initiatives to establish ISAC-enabled infrastructure for low-altitude networks [^1] https://english.www.gov.cn/news/202410/23/content_WS6718e264c6d0868f4e8ec381.html. Despite these advances, practical deployments face challenges such as dense connectivity, interference, and spectrum conflicts. To address these, joint optimization methods have been explored, including beamforming, dynamic spectrum allocation, and coordinated frequency management. For instance, [lyu2022joint] proposed a joint UAV maneuver and beamforming strategy that improved trajectory planning, bandwidth allocation, and overall energy efficiency. These optimization techniques significantly enhance system performance, as demonstrated in [10529184] and Fig. 14.
    VII Conclusion and Future Directions
    This paper has extensively examined the impact of ISAC across key domains, including operating spectrum, network architecture, sensing methodologies, security challenges, and standardization efforts within the context of 5G and beyond. These developments provide insights into ISAC’s potential and the opportunities it offers across multiple applications.
    ISAC-driven advancements have already demonstrated substantial benefits across diverse domains. In intelligent transportation, it delivers real-time road condition data to enhance autonomous driving safety and efficiency. In smart cities, it supports adaptive infrastructure that dynamically responds to residents’ needs. In industrial IoT, it facilitates high-precision monitoring and predictive maintenance, improving operational efficiency and reducing failures. These innovations not only improve existing systems but also enable entirely new services that were previously unattainable.
    As ISAC matures, its role in wireless communication will likely expand, shaping more efficient, intelligent, and adaptive networks. By embedding sensing functions into the communication infrastructure, ISAC provides wireless systems with the ability to understand and respond to the environment in real time. However, this evolution has also brought many challenges, including complex resource management mechanisms, higher-intensity security requirements, and scalable hardware integration schemes.
    Building on current progress, several open research directions can be identified:
    • Unified channel modeling: Unlike traditional communication models, ISAC demands hybrid geometric-stochastic representations that take into account multipath reflections, clutter, and scattering caused by objects. Although the continuous research of 3GPP RAN1 [R1-240] has taken an important step, future standards still need to adopt multi-bounce, clutter aware and object-tagged models to support sub-THz and mmWave sensing with high fidelity.
    • Resource optimization under multi-objective constraints: Emerging technologies such as rate-splitting multiple access (RSMA) [xu2021rate] and OTFS waveform design [gaudio2020effectiveness] show good prospects, but the existing scheduling framework rarely incorporates KPIs specific to sensing, such as angle resolution or Doppler sensitivity. Developing ISAC-aware scheduling policies that dynamically balance spatial, temporal, and spectral resources represents a compelling direction for ongoing research.
    • Privacy and Security: The exposure of environmental and user-related data introduces significant risks. Beyond conventional encryption, sensing-aware systems require native security features, such as federated access control, encrypted beamforming, and zero-trust architectures [wei2022toward]. At the same time, these mechanisms must align with evolving global data protection regulations to ensure lawful and ethical deployment.
    In conclusion, ISAC is poised to become a cornerstone of future wireless networks, unifying spectrum utilization, network intelligence, and sensing-awareness into a coherent design paradigm. Its successful realization will depend on advancing unified models, resource-aware architectures, and privacy-preserving frameworks. Ongoing interdisciplinary research will be indispensable for translating ISAC’s conceptual promise into scalable, real-world deployments, ultimately shaping the foundations of 6G and beyond. The true measure of ISAC’s success will lie not only in its technical feasibility but also in its ability to foster new applications, enhance resilience, and generate societal value across next-generation wireless ecosystems.